OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 5x = –2

hero (hero):

rewrite it as x^2 +5x + 2 = 0

hero (hero):

a = 1 b = 5 c = 2 Plug that into the formula and solve

hero (hero):

By the way, the formula is \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[x^2 + 5x = -2 \] should be rearranged such the Right Hand Side (RHS) is 0. You can do that by adding 2 to both sides of the equation.

OpenStudy (anonymous):

That would make it \[x^2 + 5x + 2 = 0\] the quadratic formular is \[(-b \pm \sqrt{b^2 - 4ac})/2a\] a is the coeffecient of x^2 so a = 1 b is the coeffecient of x so b = 5 c is the constant term so c = 2

OpenStudy (anonymous):

So my answer would be -5sqrt17/2?

OpenStudy (anonymous):

\[-5\pm \sqrt{17}\]

OpenStudy (anonymous):

Where would the 2 go?

OpenStudy (anonymous):

You mean the 2 that was on the RHS ? It would come over to the LHS and become the constant term

OpenStudy (anonymous):

Oh sorry. My mistake. I see what you mean \[(-5\pm \sqrt{17})/2\]

OpenStudy (anonymous):

I hope I got it right this time :)

OpenStudy (anonymous):

Okay thank you so much :)